what is impulse momemtum equation


Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
12 years ago

Dear student,

Impulse = Change in Momentum". In mathematical symbols, it can be written as:

Impulse = delta p

or:

Ft = mv - mv sub o = p - p sub o

or:

Ft = delta p = m delta v

ujwal reddy
12 years ago
THE DEFINITIONOF IMPULSE IS THE PRODUCT OF FORCE APPLIED AND THE TIME TAKEN I=F*T BUT F=ma if the initial velocity pf the body is u and the final velocity of the body is v then acceleration a=v-u/t therefore F=m(v-u)/t therefore substituting in the impulse equation I=mv-mu I=m(v-u) THEREFORE IMPULSE IS NOTHING BUT CHANGE IN THE MOMENTUM OF THE BODY.................!!